# WindGP: Efficient Graph Partitioning on Heterogenous Machines

### 01 Mar 2024 v1

@@ 1. Introduction
Nowadays, many data can be modeled as graphs, such as financial transactions, social network, and road network. With the strong representativity of graphs, researchers as well as industries can find out a lot valuable information by applying graph analysis algorithms (e.g., triangle counting (Zeng et al., 2022a), PageRank (Pretto, 2002) and Graph Neural Network (Hamilton et al., 2017; Wang et al., 2019; Shi et al., 2021)) on these graph data. In big-data era, distributed graph computing is widely used to process massive data in academic and industry (e.g., Huawei, Facebook, Google), supporting many important real-life applications such as financial fraud detection (Deutsch et al., 2019), social network analysis (Meng et al., 2018), and online commodity recommendation (Fan et al., 2021). Generally, the routine of distributed graph computing includes three parts: computation, communication, synchronization. This is the BSP (Bulk Synchronous Parallel) paradigm (Gerbessiotis and Valiant, 1994), which is widely used in many frameworks such as Gemini (Zhu et al., 2016) and PowerGraph (Gonzalez et al., 2012). Figure 1 shows that BSP divides the whole process into several sequential supersteps, where a barrier exist between two supersteps. In each superstep, on each machine, the communication occurs after the local computation. In the end of each superstep, all machines are synchronized to ensure that they maintain the correct information before next superstep. Thus, all machines need to wait for the slowest one, causing the long-tail effect (Lu et al., 2014). As (Lu et al., 2014) shows, with the growing data size and data complexity, the performance of distributed computing on billion-scale graphs is far from enough.
@@
With the expiration of Moore Law (Edwards, 2021), heterogeneous computing is favored by more and more organizations because various heterogeneous hardwares have mushroomed all over the area. On the one hand, machines with different configurations (e.g., CPU cores/frequency, memory capacity, network bandwidth) exist due to the evolution of hardwares (Mayer et al., 2016; Verma and Zeng, 2005). In all data centers, machines are purchased in different decades with different configurations. On the other hand, heterogeneous computing is the trend in super computing area and real-life applications due to better performance and cheaper price. In Telecom field, the machine resource in a region of carrier companies is limited and heterogeneous, where graph algorithms on billions of edges need to be performed locally to search reachable paths or find the cause of network faults. Note that telecom data can not be transfered to cloud computing due to strict data privacy. As a result, distributed graph computing on heterogeneous machines (Sulaiman et al., 2021; Masood et al., 2015; Xue et al., 2015) is becoming more and more important nowadays. This implies higher requirement and new challenges on graph partition methods.
@@
  • We design a general framework of graph partitioning, which can support fast and high-quality edge partitioning on heterogeneous machines.
@@
  • Extensive experiments on both general metric and real distributed graph algorithms show that WindGP outperforms all state-of-the-art partition methods by 1.35$\times$×$\sim$27$\times$.
@@ 2. Background
In this section, we first present the formal definition of our problem, then list the related work.
@@ 2.1. Problem Definition
This paper aims to provide fast and high-quality solutions for the problem of heterogeneous-machine graph partition (Definition 4). Obviously, this problem is edge-centric, i.e., any edge of $G$ can only exist in a single partition, but this does not hold for vertex. The details of machine parameter setting ($M_{i}$, $C_{i}^{node}$, $C_{i}^{edge}$, $C_{i}^{com}$) are in Section 5. For example, in vertex-centric and edge-centric computing systems, $C_{i}^{edge}$ and $C_{i}^{node}$ can be set to 0, respectively. As heterogeneous machines are rather different from homogeneous configurations, we propose a new metric $TC$ to measure the total time cost of each kind of edge partition. $TC$ considers both computing cost and communication cost, while the traditional replication factor (Xie et al., 2014; Zhang et al., 2017, 2021) ($RF=\frac{\sum_{u\in G}{|S(u)|}}{|V(G)|}$) only considers the communication cost ($S(u)$ represents the set of partitions that $u$ exists).
Equivalence of metrics. The new metric $TC$ theoretically corresponds to the load balance and the $RF$ metric. In homogeneous cases, $T_{i}^{cal}$ is proportional to $|V_{i}|$ or $|E_{i}|$ according to the design of distributed framework. Besides, assuming $C_{i}^{com}=1\ \forall i$ and $|V(G)|$ is fixed, we can deduce $\sum_{i}T_{i}^{com}=\sum_{u\in G}{|S(u)|\times(|S(u)|-1)}=\Theta(RF^{2})$, which is consistent with $RF$ and both of them can be used to measure the communication cost. Comparative experiments on both dense algorithms (PageRank (Pretto, 2002), Triangle (Zeng et al., 2022a)) and sparse algorithms (BFS (Buluc and Madduri, 2011), SSSP (Forster and Nanongkai, 2017)) shows that $TC$ fits the real running time with ¡10% error (see Table 1). Furthermore, given a specific graph algorithm, the computation cost of a node or an edge is the same for different nodes/edges. Therefore, $TC$ is used as the metric of partition quality in heterogeneous environments.
@@
Quantification of Machine Resource. Similar to (Zhong et al., 2023), the quantification of $M_{i},C_{i}^{node},C_{i}^{edge},C_{i}^{com}$ is done by atomic tests:
@@
HDRF 2.7G 2.3K 0.7K 1K 0.11K
NE 5.6G 4.4K 1.5K 1.8K 0.2K
    The selected graph is TW in the experiment section.
@@ 2.2. Related Work
Heterogeneous Distributed Computing. Distributed computing on big data has been well studied (Mahmud et al., 2020; C et al., 2021), but in graph computing we need to utilize the characteristics of both graphs and clusters. Recently, the heterogeneous computing on deep neural network (DNN) arises (Liu et al., 2021; Jia et al., 2022), which leads the trend of enabling heterogeneous resources. The literature includes two works for heterogeneous graph algorithms: (Verma and Zeng, 2005) only consider different computing power, while GrapH (Mayer et al., 2016) targets at various communication cost. (Verma and Zeng, 2005) coarsen $G$ first, then partition it and finally project it back to $G$. GrapH uses streaming partition and group machines into different clusters according to their network traffic. Due to the lack of collaborative optimization on machines with heterogeneous memory, computing power and network traffic, none of them can achieve a good balance between computation and communication. They use different framework and can not be combined with other solutions directly, thus we compare WindGP with them in Section 5.4. HaSGP (Zhong et al., 2023) considers the heterogenity of both computing and communication, but it has three limitations: (1) neglects different memory capacity; (2) as a streaming algorithm, lacks optimization in subgraph locality; (3) targets at high-bandwidth network and considers multi-core conflict. In its experiments, only four and 32 nodes are used and the heterogenity of both computing and communication are not considered in the same cluster. Besides, heterogeneous computing is mainly used in low-end scenarios whose network bandwidth is far smaller than memory bandwidth, thus limiting the application of HaSGP. The state-of-the-art HAEP (Zhang et al., 2023) adopts the same metrics (balance ratio $\alpha^{\prime}$ and replication factor $RF$) as homogeneous cases, and proposes heuristic neigbor expansion to improve subgraph locality. However, $\alpha^{\prime}$ and $RF$ can not depict the quality of heterogeneous partition well, as analyzed in Section 2.1. In addition, HAEP includes both computing and communication heterogenity in the same cluster, but still omits the memory heterogenity. Furthermore, the maximum cluster and the maximum dataset only contain 32 nodes and 117M edges respectively, which is not enough to prove the scalability.
@@ 5. Experiments
We evaluate our method WindGP as well as all counterparts on a machine running CentOS 7 system with 24 Intel E5-2690 2.60GHz cores, 128GB memory and 1TB disk. WindGP is implemented with C++ and compiled with g++ 7.3.0, and the optimization flag is set to -O2. According to the experiments of previous work (Zhang et al., 2017, 2021), we select four state-of-the-art algorithms as counterparts: METIS (Karypis and Kumar, 1998), HDRF (Petroni et al., 2015), NE (Zhang et al., 2017), EBV (Zhang et al., 2021). To provide a fair comparison, we modify these algorithms to meet the requirement of heterogeneous-machine edge partition, i.e., adding constraints of memory capacity of each machine. Note that METIS is originally a vertex-centric method, thus we transform it into the edge-centric solution in the same way discussed in (Zhang et al., 2017): with the node degree as the node weight, it partitions $G$ using $gpmetis$ command, then assigns each edge $\overline{uv}$ to the machine that $u$ or $v$ exists randomly as long as the machine has enough memory. In our experiments, the main metric is the $TC$ score (see Definition 4), which measures the quality of partition results. Besides, the running effect of distributed graph algorithms on different partition results is also evaluated in Section 5.4. Due to space limit, the comparison with heterogeneous solutions (Verma and Zeng, 2005; Mayer et al., 2016) is only included in the final part.
@@ 5.1. Dataset
We mainly evaluate all solutions on some representative dataset in the Stanford Network Analysis Project (SNAP (Leskovec and Krevl, 2014)) such as TW (Twitter (Kwak et al., 2010)), CO (com-Orkut), LJ(soc-LiveJournal), PO (soc-Pokec), CP (cit-Patents (Leskovec et al., 2005)) and RN (roadNet-CA). The details of these datasets are listed in Table 3. Besides, we also use a synthetic generator (R-MAT (Chakrabarti et al., 2004)) to generate a series of power-law graphs for the scalability test in Section 5.3. In order to eliminate the influence of randomness in some algorithms (e.g., NE (Zhang et al., 2017)), all tasks are run 10 times and the averaged result is recorded as the metric. In our experiments, the executing time of all partition algorithms is required to be smaller than 1 hour (for Twitter) and 10 minutes (for other datasets), respectively.
Machine Configuration. Refer to cloud platform (Garraghan et al., 2013; Herbst et al., 2018; Zakarya and Gillam, 2019), the resources of machines can be quantified by relative rates. For example, machine $i$ has the smallest memory (100GB), then we can set $M_{i}=1$ and set $M_{j}=2$ if machine $j$ has 200GB memory. It is similar for the computing ($cores\times frequency$) and communication (network bandwidth). Experiments are done on different homogeneous clusters to verify the feasibility of our methodology of resource quantification.
@@
We mainly use two types of machines: super machine and normal machine. According to most graph systems (Chen et al., 2018; Gonzalez et al., 2012; Zhu et al., 2016), $M^{node}$ and $M^{edge}$ are set to 1 and 2 respectively. Recall that in Definition 4, the configuration of each machine is a quadruple: $(M_{i},C_{i}^{node},C_{i}^{edge},C_{i}^{com})$. On large graphs (Twitter and com-Orkut), there are 100 machines (20 super machines and 80 normal machines), and the configuration of super machine and normal machine is $(10^{8},10,15,15)$ and $(3\times 10^{7},5,10,10)$, respectively. In contrast, there are 30 machines (10 super machines and 20 normal machines) on other datasets and the configuration is $(10^{7},10,15,15)$ and $(3\times 10^{6},5,10,10)$. In Section 5.3, we also study the impact of the number of machine types and the machine number.
@@
Parameter Tuning. Several hyper-parameters are used in our solution as well as others. For each algorithm, comprehensive tests are conducted on all datasets and the best parameter is found out on each dataset (details omitted). In WindGP, $\alpha$ and $\beta$ are both set to 0.3, and other parameters like $\gamma$ and $\theta$ keep the default value as specified in Section 3.

### 06 Mar 2024 v2

@@ 1. Introduction
Nowadays, many data can be modeled as graphs, such as financial transactions, social network, and road network. With the strong representativity of graphs, researchers as well as industries can find out a lot valuable information by applying graph analysis algorithms (e.g., triangle counting (Zeng et al., 2022a), PageRank (Pretto, 2002), subgraph matching (Zeng et al., 2020b; Zeng and Zou, 2018; Zeng et al., 2020a), path finding (Zeng et al., 2023) and Graph Neural Network (Hamilton et al., 2017; Wang et al., 2019; Shi et al., 2021)) on these graph data. In big-data era, distributed graph computing is widely used to process massive data in academic and industry (e.g., Huawei, Facebook, Google), supporting many important real-life applications such as financial fraud detection (Deutsch et al., 2019), social network analysis (Meng et al., 2018), and online commodity recommendation (Fan et al., 2021). Generally, the routine of distributed graph computing includes three parts: computation, communication, synchronization. This is the BSP (Bulk Synchronous Parallel) paradigm (Gerbessiotis and Valiant, 1994), which is widely used in many frameworks such as Gemini (Zhu et al., 2016) and PowerGraph (Gonzalez et al., 2012). Figure 1 shows that BSP divides the whole process into several sequential supersteps, where a barrier exist between two supersteps. In each superstep, on each machine, the communication occurs after the local computation. In the end of each superstep, all machines are synchronized to ensure that they maintain the correct information before next superstep. Thus, all machines need to wait for the slowest one, causing the long-tail effect (Lu et al., 2014). As (Lu et al., 2014) shows, with the growing data size and data complexity, the performance of distributed computing on billion-scale graphs is far from enough.
@@
With the expiration of Moore Law (Edwards, 2021), heterogeneous computing is favored by more and more organizations because various heterogeneous hardwares have mushroomed all over the area. On the one hand, machines with different configurations (e.g., CPU cores/frequency, memory capacity, network bandwidth) exist due to the evolution of hardwares (Mayer et al., 2016; Verma and Zeng, 2005). In all data centers, machines are purchased in different decades with different configurations. On the other hand, heterogeneous computing is the trend in super computing area and real-life applications due to better performance and cheaper price. In Telecom field, the machine resource in a region of carrier companies is limited and heterogeneous, where graph algorithms on billions of edges need to be performed locally to search reachable paths or find the cause of network faults. Note that telecom data can not be transfered to cloud computing due to strict data privacy, thus they can only be processed by several low-memory edge servers or even personal computers. As a result, distributed graph computing on heterogeneous machines (Sulaiman et al., 2021; Masood et al., 2015; Xue et al., 2015) is becoming more and more important nowadays. This implies higher requirement and new challenges on graph partition methods.
@@
  • We design a scalable framework of graph partitioning, which can support fast and high-quality edge partitioning on heterogeneous machines.
@@
  • Extensive experiments on both dense and sparse distributed graph algorithms show that WindGP outperforms all state-of-the-art partition methods by 1.35$\times$×$\sim$27$\times$.
@@ 2. Background
In this section, we first present the formal definition of our problem, then list the related work. The common notations within the paper are summarized in Table 2.
@@ 2.1. Problem Definition
This paper aims to provide fast and high-quality solutions for the problem of heterogeneous-machine graph partition (Definition 4). Obviously, this problem is edge-centric, i.e., any edge of $G$ can only exist in a single partition, but this does not hold for vertex. As heterogeneous machines are rather different from homogeneous configurations, we propose a new metric $TC$ to measure the total time cost of each kind of edge partition. $TC$ considers both computing cost and communication cost, while the traditional replication factor (Xie et al., 2014; Zhang et al., 2017, 2021) ($RF=\frac{\sum_{u\in G}{|S(u)|}}{|V(G)|}$) only considers the communication cost ($S(u)$ represents the set of partitions that $u$ exists).
Equivalence of metrics. The new metric $TC$ theoretically corresponds to the load balance and the $RF$ metric. In homogeneous cases, $T_{i}^{cal}$ is proportional to $|V_{i}|$ or $|E_{i}|$ according to the design of distributed framework. Besides, assuming $C_{i}^{com}=1\ \forall i$ and $|V(G)|$ is fixed, we can deduce $\sum_{i}T_{i}^{com}=\sum_{u\in G}{|S(u)|\times(|S(u)|-1)}=\Theta(RF^{2})$, which is consistent with $RF$ and both of them can be used to measure the communication cost. Comparative experiments on both dense algorithms (PageRank (Pretto, 2002), Triangle (Zeng et al., 2022a)) and sparse algorithms (BFS (Buluc and Madduri, 2011), SSSP (Forster and Nanongkai, 2017)) shows that $TC$ is proportional to the distributed running time with ¡10% error (see Table 1). Furthermore, given a specific graph algorithm, the computation cost of a node or an edge is the same for different nodes/edges. Therefore, $TC$ is used as the metric of partition quality in heterogeneous environments.
@@
Quantification of Machine Resource. Similar to (Zhong et al., 2023; Garraghan et al., 2013; Herbst et al., 2018; Zakarya and Gillam, 2019), the resources of machines can be quantified by relative rates:
@@
HDRF (Petroni et al., 2015) 2.7G 2.3K 0.7K 1K 0.11K
NE (Zhang et al., 2017) 5.6G 4.4K 1.5K 1.8K 0.2K
    The selected graph is TW on 9-machine cluster in the experiment section.
@@ 2.2. Related Work
Heterogeneous Distributed Computing. Distributed computing on big data has been well studied (Mahmud et al., 2020; C et al., 2021), but in graph computing we need to utilize the characteristics of both graphs and clusters. Recently, the heterogeneous computing on deep neural network (DNN) arises (Liu et al., 2021; Jia et al., 2022), which leads the trend of enabling heterogeneous resources. The literature includes two works for heterogeneous graph algorithms: (Verma and Zeng, 2005) only consider different computing power, while GrapH (Mayer et al., 2016) targets at various communication cost. (Verma and Zeng, 2005) coarsen $G$ first, then partition it and finally project it back to $G$. GrapH uses streaming partition and group machines into different clusters according to their network traffic. Due to the lack of collaborative optimization on machines with heterogeneous memory, computing power and network traffic, none of them can achieve a good balance between computation and communication. They use different framework and can not be combined with other solutions directly, thus we compare WindGP with them in Section 5.4. HaSGP (Zhong et al., 2023) considers the heterogeneity of both computing and communication, but it has three limitations: (1) neglects different memory capacity; (2) as a streaming algorithm, lacks optimization in subgraph locality; (3) targets at high-bandwidth network and considers multi-core conflict. In its experiments, only four and 32 nodes are used and the heterogeneity of both computing and communication are not considered in the same cluster. Besides, heterogeneous computing is mainly used in low-end scenarios whose network bandwidth is far smaller than memory bandwidth, thus limiting the application of HaSGP. The state-of-the-art HAEP (Zhang et al., 2023) adopts the same metrics (balance ratio $\alpha^{\prime}$ and replication factor $RF$) as homogeneous cases, and proposes heuristic neighbor expansion to improve subgraph locality. However, $\alpha^{\prime}$ and $RF$ can not depict the quality of heterogeneous partition well, as analyzed in Section 2.1. In addition, HAEP includes both computing and communication heterogeneity in the same cluster, but still omits the memory heterogeneity. Furthermore, the maximum cluster and the maximum dataset only contain 32 nodes and 117M edges respectively, which is not enough to prove the scalability.
@@
$num(L)$
the number of currently valid elements in set $L$
$\delta_{i}$
the precomputed capacity of partition $G_{i}$
$I_{B}(v)$
the indicator function: whether $v$ exists in other partition
$\alpha^{\prime},RF$
the load balance ratio and the replication factor in homogeneous partitioning, respectively
$\alpha$
the threshold controlling the balance between $N(u)\setminus S$ and $N(u)\cap S$
$\beta$
the threshold controlling the impact of border vertices
$\gamma$
the threshold of edges above which partitions should be destroyed
$\theta$
the proportion of edges to be removed in a partition
$N_{0},T_{0}$
the number of local and global try times
@@ 5. Experiments
We evaluate our method WindGP as well as all counterparts on a machine running CentOS 7 system with 24 Intel E5-2690 2.60GHz cores, 128GB memory and 1TB disk. WindGP is implemented with C++ and compiled with g++ 7.3.0, and the optimization flag is set to -O2. According to the experiments of previous work (Zhang et al., 2017, 2021), we select four state-of-the-art algorithms as counterparts: METIS (Karypis and Kumar, 1998), HDRF (Petroni et al., 2015), NE (Zhang et al., 2017), EBV (Zhang et al., 2021). To provide a fair comparison, we modify these algorithms to meet the requirement of heterogeneous-machine edge partition, i.e., adding constraints of memory capacity of each machine. Note that METIS is originally a vertex-centric method, thus we transform it into the edge-centric solution in the same way discussed in (Zhang et al., 2017): with the node degree as the node weight, it partitions $G$ using $gpmetis$ command, then assigns each edge $\overline{uv}$ to the machine that $u$ or $v$ exists randomly as long as the machine has enough memory. In our experiments, the main metric is the $TC$ score (Definition 4), which measures the quality of partition results. Besides, the running time of distributed graph algorithms on different partition results is also evaluated in Section 5.4. The comparison with heterogeneous solutions (Verma and Zeng, 2005; Mayer et al., 2016; Zhong et al., 2023; Zhang et al., 2023) is included in the final part.
@@ 5.1. Dataset
We mainly evaluate all solutions on some representative dataset in the Stanford Network Analysis Project (SNAP (Leskovec and Krevl, 2014)) such as TW (Twitter (Kwak et al., 2010)), CO (com-Orkut), LJ(soc-LiveJournal), PO (soc-Pokec), CP (cit-Patents (Leskovec et al., 2005)) and RN (roadNet-CA). The details of these datasets are listed in Table 3. Besides, we also use a synthetic generator (R-MAT (Chakrabarti et al., 2004)) to generate a series of power-law graphs for the scalability test in Section 5.3. In Section 5.4, graphs with billions of edges are used for distributed graph computing because modern graph and machine resource are usually very large.
To eliminate the influence of randomness in some algorithms (e.g., NE (Zhang et al., 2017)), all tasks are run 10 times and the averaged result is recorded as the metric. In our experiments, the executing time of all partition algorithms is required to be smaller than 1 hour (for graphs with billions of edges) and 10 minutes (for other datasets), respectively.
Machine Configuration. The quantification of machine resource is detailed in Section 2.1, and experiments are done on different homogeneous clusters to verify the feasibility of our methodology.
@@
We mainly use two types of machines: super machine and normal machine. For simplicity, $M^{node}$ and $M^{edge}$ are set to 1 and 2 respectively. Recall that in Definition 4, the configuration of each machine is a quadruple: $(M_{i},C_{i}^{node},C_{i}^{edge},C_{i}^{com})$. On large graphs (e.g., TW and CO), there are 100 machines (20 super machines and 80 normal machines), and the configuration of super machine and normal machine is $(10^{8},10,15,15)$ and $(3\times 10^{7},5,10,10)$, respectively. In contrast, there are 30 machines (10 super machines and 20 normal machines) on other datasets and the configuration is $(10^{7},10,15,15)$ and $(3\times 10^{6},5,10,10)$. In Section 5.3, we also study the impact of the number of machine types and the machine number.
@@
Parameter Tuning. Several hyper-parameters are used in our solution as well as others. For each algorithm, comprehensive tests are conducted on all datasets and the best parameter is found out for each dataset. In WindGP, $\alpha$ and $\beta$ are both set to 0.3, and other parameters like $\gamma$ and $\theta$ keep the default value as specified in Section 3. The details of hyper-parameter tuning in WindGP is listed in Table 4, 5, 6, 7, 8, and 9. Lookup in Table 2 for the meaning of these hyper-parameters.
As for $\alpha$, it controls the balance between $N(u)\setminus S$ and $N(u)\cap S$. On LJ, CP and RN, the average degree is low (¡7), thus the long-tail effect caused by small $N(u)\cap S$ is more prominent. While on other high-degree graphs, much larger $N(u)\setminus S$ generates too many borders which raises the communication cost. Overall, the best value is set to 0.3 for $\alpha$.
Table 4. Tuning of $\alpha$ in WindGP
$TC$ 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
TW 64M 62M 61M 60M 61M 62M 65M 70M 76M 85M
CO 34M 32M 31M 31M 32M 33M 36M 39M 45M 52M
LJ 34M 28M 25M 23M 23M 24M 25M 25M 27M 28M
PO 25M 23M 21M 21M 21M 23M 25M 29M 33M 38M
CP 20M 15M 12M 11M 11M 11M 12M 13M 13M 14M
RN 26M 20M 17M 15M 15M 16M 18M 18M 19M 19M
As for $\beta$, it controls the impact of border vertices. On graphs with low average degree, border vertices are much fewer, thus the impact of border vertex number and $N(u)\setminus S$ are both slight. In contrast, there are too many borders on TW, CO and PO, raising $TC$ in two directions. Overall, the best value is set to 0.3 for $\beta$.
Table 5. Tuning of $\beta$ in WindGP
$TC$ 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
TW 80M 71M 65M 60M 61M 62M 64M 64M 67M 70M
CO 42M 37M 33M 31M 32M 33M 33M 33M 34M 34M
LJ 24M 24M 23M 23M 24M 25M 25M 25M 26M 27M
PO 30M 25M 22M 21M 22M 22M 24M 25M 25M 26M
CP 13M 13M 12M 11M 11M 12M 12M 13M 14M 15M
RN 15M 15M 15M 15M 15M 15M 16M 17M 17M 17M
According to Algorithm 5 in Section 3.4, $\gamma$ is the threshold of edges above which partitions should be destroyed. When $\gamma$ is set to 1, only partitions with the maximum cost will be destroyed, which limits the improvement of subgraph-local search. Once $\gamma$ decreases, $TC$ drops rather slightly with much higher executing time. For example, the best $TC$ is achieved by $\gamma=0$, but the speedup is ¡9%, while the executing time of post-processing is ¿10 times longer. Besides, as the subgraph-local search needs several iterations, only a small percentage of partitions should be destroyed in each iteration. Thus, the best value is set to 0.9 for $\gamma$.
Table 6. Tuning of $\gamma$ in WindGP
$TC$ 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
TW 58M 58M 58M 58M 59M 59M 60M 60M 60M 60M 68M
CO 30M 30M 30M 30M 31M 31M 31M 31M 31M 31M 33M
LJ 22M 22M 22M 22M 22M 23M 23M 23M 23M 23M 25M
PO 20M 20M 20M 20M 21M 21M 21M 21M 21M 21M 23M
CP 10M 10M 11M 11M 11M 11M 11M 11M 11M 11M 12M
RN 15M 15M 15M 15M 15M 15M 15M 15M 15M 15M 16M
To select a appropriate proportion of edges in each destroyed machine, we vary $\gamma$ from 0.002 to 0.02 and evaluate $TC$. Obviously, too small $\gamma$ limits the optimization space of destroy-and-repair, while $\gamma$ larger than 0.1 does not bring extra gain. In conclusion, $\theta$ is set to 1% by default.
Table 7. Tuning of $\theta$ in WindGP
$TC$ 0.002 0.004 0.006 0.008 0.01 0.012 0.014 0.016 0.018 0.02
TW 67M 65M 64M 63M 60M 60M 60M 60M 60M 60M
CO 40M 38M 35M 33M 31M 31M 31M 31M 31M 31M
LJ 25M 25M 24M 24M 23M 23M 23M 23M 23M 23M
PO 26M 24M 23M 23M 21M 21M 21M 21M 21M 21M
CP 12M 12M 12M 11M 11M 11M 11M 11M 11M 11M
RN 16M 15M 15M 15M 15M 15M 15M 15M 15M 15M
As for the number of local and global try times $N_{0}$ and $T_{0}$, a thorough study is conducted that varies them from 1 to 9 for each dataset. To sum up, it is enough to set $N_{0}=5$, and the best value of $T_{0}$ varies on different graphs.
Table 8. Tuning of $N_{0}$ in WindGP
$TC$ 1 2 3 4 5 6 7 8 9
TW 68M 64M 62M 61M 60M 60M 60M 60M 60M
CO 35M 34M 32M 32M 31M 31M 31M 31M 31M
LJ 25M 24M 23M 23M 23M 23M 23M 23M 23M
PO 25M 24M 22M 22M 21M 21M 21M 21M 21M
CP 12M 12M 11M 11M 11M 11M 11M 11M 11M
RN 16M 15M 15M 15M 15M 15M 15M 15M 15M
Table 9. Tuning of $T_{0}$ in WindGP
$TC$ 1 2 3 4 5 6 7 8 9
TW 70M 66M 64M 63M 61M 61M 60M 60M 60M
CO 35M 33M 32M 32M 32M 31M 31M 31M 31M
LJ 26M 24M 24M 23M 23M 23M 23M 23M 23M
PO 25M 23M 22M 22M 21M 21M 21M 21M 21M
CP 12M 11M 11M 11M 11M 11M 11M 11M 11M
RN 17M 15M 15M 15M 15M 15M 15M 15M 15M